The most common isotope of gold is gold-197 with a measured mass of 196.966569u.

  1. Determine the mass defect;
  2. Determine the binding energy;
  3. Determine the binding energy per nucleon.
1
Expert's answer
2021-06-03T09:16:31-0400

1.

m(proton) = 1.007276 u

m(neutron) = 1.008665 u

Cu-63 has 79 protons and 118 neutrons.

m(defect) = -(196.966569 - 1.007276"\times"79 - 1.008665"\times"118) = 1.630705 u

2.

1.630705 u = 2.707849"\times"10-27 kg

E(per atom) = mc2 = 2.707849"\times"10-27 "\times" 299,792,4582 = 2.433693"\times"10-10 J

3.

E(per nuclon) = "\frac{2.433693×10^{-10} J }{197} = 1.235377\times 10^{-12} J"